# Holoported Characters: Real-time Free-viewpoint Rendering of Humans from Sparse RGB Cameras

### v1

@@ A.3 Implementation Details
Table 4: Quantitative Ablations. Here, we ablate some design choices of our SRNet module in the novel pose setting on subject S1. Note that choosing our $\mathcal{L}_{\mathrm{sr}}$ formulation and keeping a shallower architecture leads to better performance. “RF” stands for receptive field.
@@
A.4 Additional Results
We also provide additional results of our method on more subjects in the novel view and novel pose setting (Fig. 14 and Fig. 15). Our method also allows for exciting applications like texture editing (Fig. 10) and placing the character in a virtual room (Fig. 11).
A.5 End-to-end Sparse Camera Demo
In Fig. 12, we additionally show the results of our method using a 3D skeletal pose recovered from four and eight cameras, respectively, for different subjects. Note that the results with four cameras demonstrates that our method can also be integrated into an end-to-end sparse camera setup.
A.6 Additional Comparisons
We also present additional qualitative comparisons in the novel view and novel pose setting to animatable representations (Fig. 16) and real-time sparse image-driven methods (Fig. 17 and Fig. 18). In Tab. 5, we compare the quality of our reconstruction to competing methods. Our approach results in quantitative improvements in both the Chamfer (CD) and Hausdorff (HD) metrics.
Novel View Novel Pose
Method CDnormal-$\downarrow$ HDnormal-$\downarrow$ CDnormal-$\downarrow$ HDnormal-$\downarrow$
DDC 11.53 11.24 15.5 15.6
HDHumans 11.52 11.21 13.7 13.5
Ours 10.4 9.3 13.1 12.9
Table 5: Quantitative Comparison on Surface Distance. We evaluate the surface tracking in novel view and novel pose setting on subject S1. Note that, thanks to the additional Chamfer penalty, our approach outperforms previous work. This is significant in enhancing the quality of our projective texturing pipeline. Error reported in mm.
A.7 Additional Ablations
We provide additional ablation results in Fig. 9 and Tab. 4. We ablate our SRNet module by replacing it with a naive loss that is the same as $\mathcal{L}_{\mathrm{ren}}$ and find that this performs quantitatively and qualitatively worse, especially in the borders. By replacing our (shallow) SRNet architecture with a deeper architecture that utilizes a UNet [32] with high receptive field leads to worse multi-view consistency, which can be seen quantitatively (Tab. 4) and qualitatively (Fig. 9).
@@ A.7 Additional Ablations
Figure 13: Large-scale Training. Here we show novel view synthesis results from DVA trained on 300 frames and trained on 18000 frames. Note that the frame visualized is part of the training poses. While DVA is, to some extent, able to capture high fidelity details when trained with 300 frames, its quality significantly deteriorates as the number of frames increases. In contrast, our method is able to maintain high fidelity when trained on a large-scale training sequence.
A.8 ENeRF and DVA Analysis
@@
We provide more details on how we compared to ENeRF [21] and DVA [31], as they are closely related to our setting. ENeRF is a general method that generates novel views from sparse source camera views and even demonstrates impressive generalizability to completely unseen scenes. However, in their setting, they utilize nearest views at test time from a dense setup to produce the target view, which violates our requirement for only a fixed number of cameras at test time. Hence, to compare with them, we retrain their method in the same setting as ours (inference using four fixed cameras). We observe their method faces multi-view inconsistency artifacts due to a lack of reliable human priors for depth and since their final color is a simple weighted combination of the source colors (see Fig. 17 and Fig. 18).
@@
DVA achieves photorealistic telepresence in real-time from sparse cameras and 3D skeletal pose. They released a version of their source code, which relies on SMPL-X [27] for mesh tracking. We used SMPL-X tracking from twenty cameras and trained DVA in the same setting as ours. We observe that DVA can do replay and novel view synthesis with high quality if trained only on small sequences (300 frames), but their results become blurry when trained on more frames (see Fig. 13). We hypothesize that this is caused by the fact that their volume primitives have to model fine-scale deformations and appearance at the same time, while the capacity of the network is too limited to model both of those aspects of the human. Additionally, we observe that the model fails to converge on loose clothing, as their volume regularizer prevents primitives from moving far away from the SMPLX initialization, which is imperative in the case of loose clothing (see Fig. 18). Ours, in contrast, deals with deformations separately in the explicit character model, which allows us to maintain appearance quality even as the number of training frames increases.
A.9 Limitations and Future Work
@@
Though our work is a clear step towards more immersive and photorealistic avatars, there are remaining challenges yet to be addressed in the future. For example, our method, similar to other prior works, does not allow to model topological changes, e.g. opening a jacket. Future work could explore layered human representations, potentially able to model such effects. Moreover, we currently require a dense studio setup to acquire the photoreal avatar. In the future, we plan to explore more lightweight setups even for training the model. Last, our projective texturing takes the learned geometry from the deformable character model as input while not being able to further refine throughout the training. We believe that differentiable projective texturing could be an interesting direction to tackle this problem.
@@
Figure 15: Results from our method in the novel pose synthesis setting for five different subjects. The results are rendered from a novel view point not seen during training. Our method results in impressive renderings, providing very realistic wrinkle patterns and high frequency details.
@@
Figure 18: Comparison to ENeRF and DVA on Loose Clothing. ENeRF [21] can handle loose clothing reasonably well, but suffers from artifacts in other body parts (like in the hands). DVA [31] fails to converge on loose-clothed subjects, leading to artifacts in the areas like the skirt. In contrast, as our character model can handle loose clothing, we produce sharp results with details preserved.
Figure 19: Qualitative results considering the contributions of the components from our method. Without texture, results fail to reproduce fine details. Without the SR module and the features as additional input further degrades the borders and the high frequency details. Finally, without 4K training, the results are less sharp compared to our method.

### v2

@@ A.4 Implementation Details
Table 5: Quantitative Ablations. Here, we ablate some design choices in our SRNet module, in the novel pose setting on subject S1. Note that choosing our $\mathcal{L}_{\mathrm{sr}}$ formulation, and keeping a shallower architecture lead to better performance.
[…]
w/o Texture 25.82 41.36 55.17 1K
w/o Features 28.37 31.30 21.05 1K
w/o Chamfer 27.83 30.35 15.76 1K
w/o SR 28.42 31.11 20.85 1K
w/o 4K 28.85 30.50 18.01 1K
Ours 28.03 28.49 13.26 1K
Ours w/o 4K 27.72 34.49 20.89 4K
Ours 27.22 33.17 15.26 4K
Table 6: Quantitative Ablations. Here, we ablate some major design choices in the novel pose setting on subject S2 (loose clothing). Note that the efficacy of our design choices translates along subjects.
@@
A.5 Additional Results
We also provide additional results for our method on more subjects in the novel view and novel pose setting (Fig. 15 and Fig. 16). Our method also allows for exciting applications like texture editing (Fig. 10) and placing the character in a virtual room (Fig. 11).
A.6 End-to-end Sparse Camera Demo
We additionally show the result of our method using a 3D skeletal pose recovered from four and eight cameras, respectively, for different subjects (Fig. 12). Note that the result with four cameras demonstrates that our method can also be integrated into an end-to-end sparse camera setup.
Appendix B Additional Comparisons
We present additional qualitative comparisons in the novel view and novel pose setting to animatable representations (Fig. 17) and real-time sparse image-driven methods (Fig. 18 and Fig. 20). We also present a zoomed-in face comparison of our method to show our efficacy in capturing face expressions (Fig. 13). In Tab. 3, we compare the quality of our geometry reconstruction against competing methods. Our approach provides quantitative improvements in the Chamfer (CD) and Hausdorff metrics (HD).
B.1 Additional Ablations
Here, we provide some additional ablation results (Fig. 9 and Tab. 5). We ablate our SRNet module by replacing it with a naive loss that is the same as $\mathcal{L}_{\mathrm{ren}}$ and find that this performs quantitatively worse and also qualitatively (especially at the borders). Also, replacing a shallow SRNet architecture with a deeper architecture that utilizes UNet [37], and additional upsampling layers leads to worse multi-view consistency, which can be seen quantitatively and qualitatively. Additionally, we also add an ablation for our major components, on a subject in loose clothing (Tab. 6).
@@ B.1 Additional Ablations
Figure 13: Comparison with a zoomed-in virtual view. Here we present a result, where we render a view close to the face. Notice the ability of our method to capture facial expressions with much higher fidelity.
@@
Figure 14: Large-scale Training. We present a novel view synthesis (replay), the result from DVA when it is trained on 300 frames, and when it’s trained on 18000 frames. Note that the frame visualized is part of the training frames. While DVA can capture high-fidelity details, when trained with fewer frames, the performance deteriorates as the number of frames increases. In contrast, ours maintains high fidelity, even as the number of frames increases.
B.2 ENeRF and DVA Analysis
@@
We additionally provide more details on how we compared to ENeRF [23] and DVA [36], as they are closely related to our setting. ENeRF is a general method that generates novel views from sparse source camera views and even demonstrates impressive generalizability to completely unseen scenes. However, in their setting, they utilize nearest views at test time from a dense setup to produce the target view, which violates our requirement for only a fixed number of cameras at test time. Hence, to compare with them, we retrain their method in the same setting as ours (inference using four fixed cameras). We observe their method faces multi-view consistency artifacts due to a lack of reliable human priors for depth, and as their final color is a weighted combination of the source colors (see Fig. 18 and Fig. 20).
@@
DVA achieves photorealistic telepresence in real-time from sparse cameras and 3D skeletal pose. They released a version of their source code, which relies on SMPLX [32] for mesh tracking and provided scripts to reproduce results on the ZJU dataset [34]. First, we reproduce results on the ZJU dataset. Then, to test their robustness to loose clothing and challenging poses, we evaluate on the DynaCap dataset. We used SMPLX tracking from twenty cameras and trained their method in the same setting as ours. We observe while they can do replay and novel view synthesis with high quality if we train only on small sequences (300 frames), their result becomes blurry when trained on more frames (see Fig. 14). We hypothesize that this is caused by the fact that their volume primitives have to model fine-scale deformations and appearance at the same time, while the capacity of the network is too limited to model both of those aspects of the human. Additionally, we observe that the model fails to converge on loose clothing, as their volume regularizer prevents primitives from moving far away from the SMPLX initialization, which is imperative in the case of loose clothing (see Fig. 20). Ours, in contrast, deals with deformations separately in the explicit character model, which allows us to maintain appearance quality even as the number of training frames increases.
B.3 Limitations and Future Work
@@
Though our work is a clear step towards more immersive and photorealistic avatars, there are remaining challenges yet to be addressed in the future. For example, our method, similar to other prior works, does not allow to model topological changes, e.g. opening a jacket. Future work could explore layered human representations, potentially able to model such effects. Flickering artifacts occur in our method, due to inconsistent color calibration of the multi-view cameras. We believe a joint optimization of camera parameters, i.e. color, extrinsic, and intrinsic calibration, can potentially resolve this limitation. Additionally, tracking errors in the case of fast motions, e.g. jumping, may result in artifacts in the renderings. Tightly entangling the tracking and rendering might resolve this in the future. Moreover, we currently require a dense studio setup to acquire the photoreal avatar. In the future, we plan to explore more lightweight setups even for training the model. Last, our projective texturing takes the learned geometry from the deformable character model as input while not being able to further refine it throughout the training. We believe differentiable projective texturing could be an interesting direction to tackle this problem.
@@
Figure 16: Results from our method in the novel pose synthesis setting for five different subjects. The results are rendered from a novel view point not seen during training. Our method results in impressive renderings, providing very realistic wrinkle patterns and high-frequency details.
@@
Figure 19: Qualitative results considering the contributions of the components from our method. Without texture, results fail to reproduce fine details. Without the SR module and the features as additional input further degrades the borders and the high-frequency details. Finally, without 4K training, the results are less sharp compared to our method.
Figure 20: Comparison to ENeRF and DVA on Loose Clothing. ENeRF [23] can handle loose clothing reasonably well, but suffers from artifacts in other body parts (like in the hands). DVA [36] fails to converge on loose-clothed subjects, leading to artifacts in areas like the skirt. In contrast, as our character model can handle loose clothing, we produce sharp results with details preserved.